OpenStudy (trojanpoem):

Question : Help please ! 0.4122 grams of an organic compound was burnt totally , so the weight of the containers of absorbing water vapor and carbon dioxide were increased by 0.3618 grams , 0.762 grams respectively. 1- Find the percentage of carbon and hydrogen in the compound. 2- Is there any other element in the organic compound ? , Giving a reason.

OpenStudy (phi):

How far did you get ?

OpenStudy (michele_laino):

from your data, we can see that the ratio: water vapor and carbon dioxide =0.3618/ 0.762= 1/2 (approximated value)

OpenStudy (trojanpoem):

What confuses me is that Organic compound -(burnt)-> nH2o + mCo2 so the sum of their masses must be equal or more to the compound itself but it is more.

OpenStudy (trojanpoem):

I meant equal or less if equal -> Only H , C if less H, C, other elements

OpenStudy (phi):

I assume the O comes from the air, and the organic compound does not contain O ?

OpenStudy (trojanpoem):

Maybe and maybe it was C2H5OH or something else with O

OpenStudy (trojanpoem):

But I think you are right Phi, the O comes from the air as 0.3618 + 0.762 (not equal to) 0.4122

OpenStudy (trojanpoem):

if so the mass of (O) = 0.3618 + 0.762 - 0.4122 = 0.7116 grams

OpenStudy (phi):

0.4122 grams of an organic compound was burnt totally I would like to know the weight of any solid ash. but you have (CH other stuff) + O= H2O + CO2 (+ ash?) so the total weight of H2O + CO2 - compound is the weight of added O ?

OpenStudy (trojanpoem):

Yeah.

OpenStudy (trojanpoem):

The problem already asks about ashes ( other elements) in the next part (2).

OpenStudy (phi):

my original thought was 2/18 of the H2O represents the H in the compound and 12/44 of the CO2 is the amount of C

OpenStudy (trojanpoem):